# Avoiding Catastrophic Forgetting in Visual Classification Using Human Concept Formation

### v1

@@ A New Tensor Representation
Baseline Neural Network Architectures
We conducted a comprehensive performance comparison between Cobweb/4V and two baseline neural network architectures. The first one employs fully connected layers (fc), a versatile choice for visual classification with demonstrated effectiveness across multiple tasks (?, ?, ?, ?). The second baseline incorporates additional convolutional neural network (CNN) layers (fc-cnn), which have proven effective for computer vision tasks (?, ?, ?).
To investigate catastrophic forgetting, we set the number of hidden layers in the fc network to 1 to be directly comparable to Cobweb/4V, which does not use a multi-layer architecture. The fc-cnn network is more representative of a typical CNN for MNIST, with ten $5\times 5$ convolutional filters, a $2\times 2$ max-pooling layer, twenty $5\times 5$ convolutional filters, another $2\times 2$ max-pool layer, and one fully connected layer. Additionally, both fc and fc-cnn networks use ReLU activation functions and are trained with 5 epochs, 0.9 SGD momentum, a hidden-layer size of 128, a batch size of 64, and a learning rate of 0.00365.
Methods and Hypotheses
@@
Experiment 1 (Learning with Modest Data)
@@
We expect each approach to improve its performance given more training examples. Notably, we expect that the neural network baselines will require more data to achieve comparable performance to cobweb4v. This is attributed to the substantial number of gradient updates needed for neural networks to converge, whereas cobweb4v operates differently. Given the extensive MNIST training dataset and the success of CNNs on this dataset in prior work (?, ?, ?), we anticipate that fc-cnn may surpass cobweb4v at some point in the training process, ultimately achieving greater performance.
Experiment 2 (Learning without Forgetting)
@@
Given that neural nets are known to exhibit forgetting, we also tested neural network variations that utilize replay to mitigate this effect. Each replay approach (labeled fc(-cnn)-replay), maintains a replay buffer with 1,000 examples. Given each new split, the network is trained with all the data in the replay buffer and all the data in the incoming split. After training, we randomly sample 1,000 examples from the buffer and split data. This sample is then carried forward in the buffer for the next training iteration. Based on prior work by ? (?), we chose to use replay because their analysis suggests it outperforms alternatives, such as regularization. [^4] We also tested EWC, but found its performance comparable to the standard NN approaches, so we do not report it here. Additionally, we employ sampling-based replay because other work suggests that rehearsing tasks on a fading schedule is particularly effective for mitigating forgetting (?, ?). Our sampling approach emulates faded practice, as all digits (even the chosen one that only appears in the first split) have some fading probability of being rehearsed at all subsequent splits.
We anticipate that fc(-cnn) will exhibit catastrophic forgetting of the chosen digit after the initial training split. As these approaches encounter subsequent training splits that exclude the chosen digit, their proficiency in classifying it is expected to rapidly decline. This phenomenon in neural networks is attributed to weight adjustments accommodating new task objectives, leading to the replacement of prior task knowledge (?, ?). We expect fc(-cnn)-replay will be more robust to forgetting, but once the chosen digit data falls out of the finite buffer, it will also exhibit catastrophic forgetting. In contrast, cobweb4v is expected to maintain consistent performance on the chosen digit, even when being repeatedly trained on new data that excludes this digit. Based on prior work (?, ?, ?), we expect it may exhibit some decrease in performance as new training data interferes with its ability to correctly categorize the chosen digit; however, the decrease in performance will not be catastrophic.
Results
@@
Figure 3 shows the results of Experiment 2, illustrating the variations in the summarized test accuracy of each approach on the chosen-digit test set after processing each of the 10 training splits. Despite all approaches learning all available digits after the first split, the accuracy of neural network baselines without replay still experiences a sharp decline and finally reaches an accuracy near zero. With replay, the neural networks have less performance decay initially, but as hypothesized, they also catastrophically forget when presented with an increasing number of splits. In contrast, the accuracy of cobweb4v gradually declines (we suspect due to interference) and maintains a respectable accuracy of 0.937 after training on 9 splits that do not contain the target digit. These results suggest cobweb is robust to catastrophic forgetting effects, outperforming all neural net approaches, even those using replay.
@@ Discussion
In Experiment 1, we examine cobweb4v’s ability to learn visual concepts with varying numbers of training samples. The results show it is data efficient, learning more quickly than the neural net methods. We also find that it achieves comparable performance to the fc approach. The fc-cnn system ultimately achieves slightly better accuracy than both cobweb4v and fc, but this is to be expected as it has an additional representation learning approach (convolutional layers) that has been shown to increase performance on this task. We hypothesize that if Cobweb also utilized a representation learning technique, such as convolutional processing, it would also achieve similar higher performance. Some preliminary efforts have started to explore this research direction (?, ?). It is worth noting that, in this experiment, the neural networks utilize a smaller batch size of 5, which was necessary given the structure of the incremental learning task. While a larger batch size might enhance performance, it may compromise generalization ability, as highlighted by ? (?). In real-world scenarios, managing the information scale in continuous learning of visual concepts poses challenges. Cobweb/4V showcases human-like learning by rapidly acquiring and refining its learned structure with limited data. This underscores the efficiency of Cobweb/4V in learning visual concepts, aligning closely with human learning behavior (?, ?). The results of Experiment 2 show that neural networks experience significant performance decay after learning the chosen digit due to catastrophic forgetting. Additionally, we observe that neural nets utilizing replay—a strategy anticipated to mitigate the forgetting effect—exhibit less initial forgetting, but eventually forget catastrophically as the number of splits increases. In contrast, cobweb4v avoids catastrophic forgetting, displaying only limited interference as it experiences more of the non-chosen digits. It consistently outperforms all neural network baselines and retains concepts effectively, highlighting Cobweb/4V’s continual learning superiority.
@@
In light of our experimental findings, the unique attributes exhibited by Cobweb/4V in the context of continual learning deserve particular attention. On one hand, Cobweb/4V dynamically adapts its structure based on the incoming data to minimize the impact of forgetting, which echoes the principles of architectural-based methods reviewed in our study. For example, Cobweb/4V approach shares some similarity with neural network approaches like cascade correlation (?, ?), hinting at their potential for mitigating catastrophic forgetting. On the other hand, Cobweb/4V shares some alignment with replay-based methods leveraging its tree structure to selectively revisit previously learned concepts that are relevant to the incoming instances. We believe Cobweb/4V’s approach might suggest new strategies for replay, such as retaining prototypes or prototypical examples (akin to Cobweb’s intermediate concept nodes) in the buffer. Our hope in the current study is that we might extract key lessons from Cobweb/4V’s approach that might generalize to other machine learning methods, such as neural networks.
Conclusion and Future Work

### v2

@@ 4.1 Preliminary Experiment
4.2 Baseline Neural Network Architectures
We conducted a comprehensive performance comparison between Cobweb/4V and two baseline neural network architectures. The first one employs fully connected layers (fc), a versatile choice for visual classification with demonstrated effectiveness across multiple tasks (Rumelhart et al., 1986; LeCun et al., 1998; Krizhevsky et al., 2009). The second baseline incorporates additional convolutional neural network (CNN) layers (fc-cnn), which have proven effective for computer vision tasks (LeCun et al., 1998; Gu et al., 2018).
To investigate catastrophic forgetting, we set the number of hidden layers in the fc network to 1 to be directly comparable to Cobweb/4V, which does not use a multi-layer architecture. The fc-cnn network is more representative of a typical CNN for MNIST, with ten $5\times 5$ convolutional filters, a $2\times 2$ max-pooling layer, twenty $5\times 5$ convolutional filters, another $2\times 2$ max-pool layer, and one fully connected layer. To tune the hyperparameters, a grid search was employed with the fc network to identify optimal values.Both fc and fc-cnn networks use ReLU activation functions and are trained with 5 epochs, 0.9 SGD momentum, a hidden-layer size of 128, a batch size of 64, and a learning rate of 0.00365.
4.3 Methods and Hypotheses
@@
4.3.1 Experiment 1 (Learning with Modest Data)
@@
We expect each approach to improve its performance given more training examples. Notably, we expect that the neural network baselines will require more data to achieve comparable performance to cobweb4v. This is attributed to the substantial number of gradient updates needed for neural networks to converge, whereas cobweb4v operates differently. Given the extensive MNIST training dataset and the success of CNNs on this dataset in prior work (LeCun et al., 1998; Krizhevsky et al., 2012), we anticipate that fc-cnn may surpass cobweb4v at some point in the training process, ultimately achieving greater performance.
4.3.2 Experiment 2 (Learning without Forgetting)
@@
Given that neural nets are known to exhibit forgetting, we also tested neural network variations that utilize replay to mitigate this effect. Each replay approach (labeled fc(-cnn)-replay), maintains a replay buffer with 1,000 examples. Given each new split, the network is trained with all the data in the replay buffer and all the data in the incoming split. After training, we randomly sample 1,000 examples from the buffer and split data. This sample is then carried forward in the buffer for the next training iteration. Based on prior work by Van de Ven & Tolias (2019), we chose to use replay because their analysis suggests it outperforms alternatives, such as regularization. [^5] We also tested EWC, but found its performance comparable to the standard NN approaches, so we do not report it here. Additionally, we employ sampling-based replay because other work suggests that rehearsing tasks on a fading schedule is particularly effective for mitigating forgetting (Gandhi et al., 2024). Our sampling approach emulates faded practice, as all digits (even the chosen one that only appears in the first split) have some fading probability of being rehearsed at all subsequent splits.
We anticipate that fc(-cnn) will exhibit catastrophic forgetting of the chosen digit after the initial training split. As these approaches encounter subsequent training splits that exclude the chosen digit, their proficiency in classifying it is expected to rapidly decline. This phenomenon in neural networks is attributed to weight adjustments accommodating new task objectives, leading to the replacement of prior task knowledge (Kirkpatrick et al., 2017). We expect fc(-cnn)-replay will be more robust to forgetting, but once the chosen digit data falls out of the finite buffer, it will also exhibit catastrophic forgetting. In contrast, cobweb4v is expected to maintain consistent performance on the chosen digit, even when being repeatedly trained on new data that excludes this digit. Based on prior work (Feigenbaum, 1961; Crowder, 2014), we expect it may exhibit some decrease in performance as new training data interferes with its ability to correctly categorize the chosen digit; however, the decrease in performance will not be catastrophic.
4.4 Results
@@
Figure 6 shows the results of Experiment 2, illustrating the variations in the summarized test accuracy of each approach on the chosen-digit test set after processing each of the 10 training splits. Despite all approaches learning all available digits after the first split, the accuracy of neural network baselines without replay still experiences a sharp decline and finally reaches an accuracy near zero. With replay, the neural networks have less performance decay initially, but as hypothesized, they also catastrophically forget when presented with an increasing number of splits. In contrast, the accuracy of cobweb4v gradually declines due to feature interference, where concepts at a branch point share similar features, which makes it difficult to determine which branch to follow during categorization. Unlike neural networks, which propagate activation over all nodes, Cobweb only activates a relatively small portion of its nodes during categorization ($N_{max}$). This is more computationally efficient, but may make it more susceptible to interference effects. The results show that cobweb4v maintains a respectable accuracy of 0.937 after training on 9 splits that do not contain the target digit, which suggests that, despite some interference, Cobweb is robust to catastrophic forgetting. Further, it outperforms all neural network approaches, even those using replay.
@@ 4.5 Discussion
In Experiment 1, we examine cobweb4v’s ability to learn visual concepts with varying numbers of training samples. The results show it is data efficient, learning more quickly than the neural net methods. We also find that it achieves comparable performance to the fc approach. The fc-cnn system ultimately achieves slightly better accuracy than both cobweb4v and fc, but this is to be expected as it has an additional representation learning approach (convolutional layers) that has been shown to increase performance on this task. This can be explained by the effectiveness of intermediate representations, which as mentioned, are more effective for generalization compared to the individual pixels. We hypothesize that if Cobweb also utilized a representation learning technique, it would also achieve similar higher performance. Some preliminary efforts have started to explore this research direction (MacLellan & Thakur, 2022). It is worth noting that, in this experiment, the neural networks utilize a smaller batch size of 5, which was necessary given the structure of the incremental learning task. While a larger batch size might enhance performance, it may compromise generalization ability, as highlighted by Keskar et al. (2016). In real-world scenarios, managing the information scale in continuous learning of visual concepts poses challenges. Cobweb/4V showcases human-like learning by rapidly acquiring and refining its learned structure with limited data. This underscores the efficiency of Cobweb/4V in learning visual concepts, aligning closely with human learning behavior (Langley, 2022).
The results of Experiment 2 show that neural networks experience significant performance decay after learning the chosen digit due to catastrophic forgetting. Additionally, we observe that neural nets utilizing replay—a strategy anticipated to mitigate the forgetting effect—exhibit less initial forgetting, but eventually forget catastrophically as the number of splits increases. In contrast, cobweb4v avoids catastrophic forgetting, displaying only limited interference as it experiences more of the non-chosen digits. It is important to note that while Cobweb is instance-based, it also uses prototypes (the intermediate concepts) to generate predictions. In fact, when $N_{max}$ is small relative to the size of the tree, then it may not use any leaf concepts (which represent specific instances) during prediction. Thus, there is no guarantee that Cobweb will be robust against forgetting. Cobweb uses intermediate concepts as an index to guide dynamic access to its tree structure during categorization. As mentioned previously, sometimes interference leads to poor categorization decisions that might prevent it from accessing the relevant portions of its tree. This provides a computational explanation for what forgetting is—a loss of access to previously available memories due to interference. Therefore, it is plausible that Cobweb may still encounter challenges related to catastrophic forgetting. Nevertheless, Cobweb/4V consistently outperforms all neural network baselines and retains concepts effectively, highlighting Cobweb/4V’s continual learning superiority.
@@
In light of our experimental findings, the unique attributes exhibited by Cobweb/4V in the context of continual learning deserve particular attention. On one hand, Cobweb/4V dynamically adapts its structure based on the incoming data to minimize the impact of forgetting, which echoes the principles of architectural-based methods reviewed in our study. For example, Cobweb/4V approach shares some similarity with neural network approaches like cascade correlation (Fahlman & Lebiere, 1989), hinting at their potential for mitigating catastrophic forgetting. On the other hand, Cobweb/4V shares some alignment with replay-based methods leveraging its tree structure to selectively revisit previously learned concepts that are relevant to the incoming instances. We believe Cobweb/4V’s approach might suggest new strategies for replay, such as retaining prototypes or prototypical examples (akin to Cobweb’s intermediate concept nodes) in the buffer. Our hope in the current study is that we might extract key lessons from Cobweb/4V’s approach that might generalize to other machine learning methods, such as neural networks.
5 Conclusion and Future Work